OpenStudy (anonymous):

What is the sum of the arithmetic sequence 9, 14, 19, …, if there are 32 terms?

OpenStudy (anonymous):

Here,a =9 d = 14 - 9= 5 and n = 32.. For finding last term: \[a _{n} = a + (n - 1)d\] Put the above values and find a_n.. Now, for sum: \[Sum = n(a + a_n)/2\] and get the desired sum... Able to do it or not??

OpenStudy (anonymous):

ya thats what i di and i got 164, but thats not a choice

OpenStudy (anonymous):

2,768 2,800 2,832 2,864

OpenStudy (anonymous):

those are the options

OpenStudy (anonymous):

What you are getting the last term??

OpenStudy (anonymous):

ohh ya nevermind, i just did the first part, ill try again

OpenStudy (anonymous):

i got 2768

OpenStudy (anonymous):

You did the first part only.. Last term is 164.. Use the second formula I have written and find the sum.. Got or not???

OpenStudy (anonymous):

You are absolutely correct dear... Welcome dear..

OpenStudy (anonymous):

thanks